• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name riordan

    Log entry time 04:19:44 on October 11, 2010

    Entry number 334380

    keyword=Stopping for tonight - hall to controlled access

    The EPICS gateway will need to be fixed before we can take useful beam. Aside from not being able to control the detector HV and look at the BPMs, we cannot read or change the magnets and the dipole cannot regulate itself (and is already significantly off according to Alex), making spectrometer calibration useless.

    After talking with Alex, we've decided to stop for the night, as we would only have an hour or two of beam once the software on call got in. I've asked for the hall to go to controlled access.

    Owl shift is on standby. Day shift tomorrow is on standby and will be called in if the injector work is finished before 4:00PM. Call the run coordinator at 876-1787 if you have questions.



    A copy of this log entry has been emailed to: ellie, mfriend